OpenStudy (anonymous):

If h(x) = 3x − 1 and j(x) = −2x, find h[j(2)]

OpenStudy (amistre64):

what is j(2) equal?

OpenStudy (amistre64):

h(j(a)) can be decomposed to: j(a) = b ; therefore h(b) = h(j(a))

OpenStudy (amistre64):

j(2) = # h(#) = h(j(2))

OpenStudy (anonymous):

If h(x) = 3x − 1 and j(x) = −2x, find h[j(2)] h[ j(x)] = 3 (-2x) - 1 = 6x - 1 -> h[j(2)] = 6 * 2 - 1 = 11